OpenStudy (rock_mit182):

Two mechanics worked on a car. The first mechanic worked for hours, and the second mechanic worked for hours. Together they charged a total of . What was the rate charged per hour by each mechanic if the sum of the two rates was per hour?

OpenStudy (anonymous):

Your question seems to be missing numbers..? Or is it just me?

OpenStudy (anonymous):

yes it is missing numbers

OpenStudy (rock_mit182):

Two mechanics worked on a car. The first mechanic worked for 10 hours, and the second mechanic worked for 15 hours. Together they charged a total of 1800 . What was the rate charged per hour by each mechanic if the sum of the two rates was 155 per hour?

OpenStudy (rock_mit182):

(1800 and 155) dolars

OpenStudy (rock_mit182):

@Ashleyisakitty @whpalmer4

OpenStudy (whpalmer4):

let's call the mechanics Fred and Sue. Fred's rate will be \(f\) and Sue's rate will be \(s\) \[f+s = 155\]because the total rate is $155/hr. Good so far?

OpenStudy (rock_mit182):

yeah sure

OpenStudy (whpalmer4):

Now Fred worked 10 hours and Sue worked 15, so the amount billed by Fred is how much?

OpenStudy (whpalmer4):

remember, \(f\) is Fred's hourly rate

OpenStudy (rock_mit182):

@whpalmer4 is it: f = 150 -s

OpenStudy (whpalmer4):

No, Fred works 10 hours, at a rate of \(f\) per hour. For 10 hours work, he bills \(10f\), right?

OpenStudy (rock_mit182):

and the other one 15s

OpenStudy (whpalmer4):

Similarly, Sue bills \(15s\) for working 15 hours. Right! So can you write an equation that shows how they get to $1800 for the total bill?

OpenStudy (rock_mit182):

so the systme would be: 10f + 15s =1800 f +s = 150

OpenStudy (whpalmer4):

Close, but not quite. Look again at the sum of the two rates...

OpenStudy (rock_mit182):

I cannot fail this question ! please

OpenStudy (whpalmer4):

Hey, if you get it wrong with my help, I'll be very surprised :-)

OpenStudy (rock_mit182):

oh yeah its is 155 in the secon equation

OpenStudy (rock_mit182):

lol yeah i know you're a guru on math

OpenStudy (whpalmer4):

with story problems my specialty :-)

OpenStudy (whpalmer4):

are you solving by substitution or elimination?

OpenStudy (whpalmer4):

would you like me to stop you now, or after you've calculated the other half of the wrong answer? :-)

OpenStudy (whpalmer4):

\[10f+15s = 1800\]\[(-10)*f+(-10)*s = (-10)*155\]\[-10f -10s = -1550\] Why don't you try using that as your second equation...

OpenStudy (rock_mit182):

i see more clear now lol

OpenStudy (whpalmer4):

you know that if Sue doesn't charge anything for her time, then Fred charges 155/hr. Any answers that have someone making more than 155 are clearly wrong :-)

OpenStudy (rock_mit182):

s = 50

OpenStudy (whpalmer4):

get the other one too

OpenStudy (whpalmer4):

what did I just say?

OpenStudy (rock_mit182):

ok again i failed for do thigns too fast

OpenStudy (rock_mit182):

100

OpenStudy (whpalmer4):

plug those into original formulas, does the answer make sense?

OpenStudy (rock_mit182):

for the first one it doesn't

OpenStudy (whpalmer4):

okay, that means it isn't correct. maybe try with substitution instead?

OpenStudy (rock_mit182):

ok

OpenStudy (whpalmer4):

how are we doing?

OpenStudy (rock_mit182):

well i was on determinants by ithinkj i failed again

OpenStudy (whpalmer4):

okay, let's go through the solution both with substitution and elimination. \[10f + 15s = 1800\]\[f+s = 155\]Agreed? Check your problem carefully, I don't want to be blamed for doing the wrong problem :-)

OpenStudy (whpalmer4):

Fred works 10 hours, Sue works 15, they bill at a combined rate of $155, and the bill is $1800.

OpenStudy (rock_mit182):

10(-s +155) +15s = 1800

OpenStudy (whpalmer4):

Okay, good... keep going

OpenStudy (rock_mit182):

5s = 250 --> 50

OpenStudy (rock_mit182):

ok something is right. s!

OpenStudy (whpalmer4):

Okay, what is Fred's rate if Sue's rate is 50?

OpenStudy (rock_mit182):

f =1050/10

OpenStudy (rock_mit182):

or 105

OpenStudy (whpalmer4):

Plug both into the original equations (both of them). Do you get true statements from those answers? Remember, it must satisfy all of the equations to be correct.

OpenStudy (rock_mit182):

yeah now it is right for the first one and the second one

OpenStudy (rock_mit182):

oh man thanks a lot, You have been very helpful, are you a professor ?

OpenStudy (rock_mit182):

10(105) + 15(50) = 1800 105 + 50 = 155 both true

OpenStudy (whpalmer4):

Okay, let's do it by elimination: \[10f + 15s = 1800\]\[f+s = 155\]I'm going to multiply the second one by -10 \[10f+15s = 1800\]\[-10f-10s=-1550\] Agreed?

OpenStudy (whpalmer4):

why don't you finish it off...

OpenStudy (rock_mit182):

ok i will

OpenStudy (rock_mit182):

0 + 5s = 250 --> 50

OpenStudy (rock_mit182):

1) 10f +15(50) = 1800 10 f = 1800 -750 f = 1050/10 --> 105

OpenStudy (whpalmer4):

I got excited earlier when you got s=50, but then it got away from you somehow. You should go back and look at what you did and try to understand what the mistake was, and why you made it. Learn from your mistakes! For example, I know that subtracting negative quantities is something that has a much higher chance of error when I do it, so I try to arrange my work to do less of that (and check it very carefully) when possible.

OpenStudy (rock_mit182):

(f,s) --> (105,50)

OpenStudy (rock_mit182):

brother seriously I'm so thankful, I'm going to verify with the determinants too cause any method has to work out

OpenStudy (whpalmer4):

ah, you know the matrix route, too? good! I don't do it often enough to be confident in my abilities...

OpenStudy (rock_mit182):

lol is long but i like it

OpenStudy (whpalmer4):

you could also do this in your head via a different route :-) assume that fred bills at the whole $155/hr. he works 10 hours, so that makes only $1550. That's $250 shy of the total bill. Now, Sue works 15 hours, so every dollar she bills adds 15 to the total, but 10 comes off the total because Fred doesn't bill that dollar, so giving a dollar/hour to Sue increases the take by a total of $5. You need to boost the take by $250, divided by $5/(dollar/hour) for Sue gives you a rate of 50/hour for Sue. Easy, right?

OpenStudy (rock_mit182):

amazing, i never thought that

OpenStudy (rock_mit182):

what if i asume the oppsite, i mean starting with sue

OpenStudy (rock_mit182):

i would have the f, right ?

OpenStudy (whpalmer4):

That works too. She bills 15 hours, so that's 15*155 = 2325 (a bit harder to compute, that's why I did the other one!). That's 2325-1800 = 525 too high. Now the same calculation applies: every $/hour taken away from Sue cuts the total by $5, so what is $525/5? 105...

OpenStudy (rock_mit182):

mmm i see,

OpenStudy (rock_mit182):

with determinants it seems is impossible but anyway

OpenStudy (whpalmer4):

if you want to really solidify your story problem skills, there's a good book called How to Solve Word Problems in Algebra by Mildred Johnson. she breaks down the dozen or so basic problem types (coins, mixtures, rates, etc.) and really does a good job of it, I think. See if you local library has a copy, or pick up a copy at the bookstore, it's an inexpensive paperback as I recall.

OpenStudy (rock_mit182):

ok bro I will search that book , maybe it is on internet

OpenStudy (rock_mit182):

see you the next time, bye